# MVP-Shot: Multi-Velocity Progressive-Alignment Framework for Few-Shot Action Recognition

### v1

@@ Abstract
Recent few-shot action recognition (FSAR) methods achieve promising performance by performing semantic matching on learned discriminative features. However, most FSAR methods focus on single-scale (e.g., frame-level, segment-level, etc.) feature alignment, which ignores that human actions with the same semantic may appear at different velocities. To this end, we develop a novel Multi-Velocity Progressive-alignment (MVP-Shot) framework to progressively learn and align semantic-related action features at multi-velocity levels. Concretely, a Multi-Velocity Feature Alignment (MVFA) module is designed to measure the similarity between features from support and query videos with different velocity scales and then merge all similarity scores in a residual fashion. To avoid the multiple velocity features deviating from the underlying motion semantic, our proposed Progressive Semantic-Tailored Interaction (PSTI) module injects velocity-tailored text information into the video feature via feature interaction on channel and temporal domains at different velocities. The above two modules compensate for each other to predict query categories more accurately under the few-shot settings. Experimental results show our method outperforms current state-of-the-art methods on multiple standard few-shot benchmarks (i.e., HMDB51, UCF101, Kinetics, and SSv2-small).
@@
Few-shot action recognition, Few-shot learning, Action recognition, Pre-trained vision-language model
@@ I Introduction
Due to the increasing demand for video analysis in real-world scenarios, action recognition has gained significant progress [1, 2, 3, 4, 5] in recent years. Despite this, these successes still heavily rely on extensive manual data annotation and collecting these data is expensive and labor-intensive. This significant bottleneck severely constrains the broader applicability of this task. To relieve dependence on a large amount of labeled data, few-shot action recognition (FSAR) is a promising direction and has gained widespread attention [6, 7, 8, 9, 10] in recent years. It aims to recognize new categories with only a few labeled examples.
@@
Current mainstream FSAR approaches follow the metric-based meta-learning paradigm, which maps videos into an appropriate feature space and performs feature alignment to predict action category labels. Most metric-based methods perform frame-level alignment [11, 7, 12, 13, 9] on learned discriminative frame-level features, which measure the similarity between query and support videos via diverse matching strategies [11, 7, 14]. In addition, some methods perform segment-level alignment [15, 16, 17] on obtained segment-level features. Typically, TARN [15] first uniformly samples segments from videos, and performs segment-level alignment based on segment features extracted by 3D-CNN. Although existing FSAR methods have achieved remarkable performance, these methods still mainly focus on single-scale feature alignment (i.e., frame-level alignment and segment-level alignment), as shown in Fig. 1(a), which is insufficient to reflect the diversity of action speeds. Therefore, these single-scale approaches often result in inaccurate matching when dealing with action instances with diverse speeds. In real-world scenarios, different individuals may perform the same action class at different velocities. Take the action “ride horse” as an example, different persons act at different velocities. Specifically from a visual perspective, certain horses may be perceived sprinting at a rapid pace, while others may be observed moving at a more leisurely walk.
@@
In our paper, we focus on multi-velocity feature alignment in FSAR, where multi-velocity features are utilized to capture pair-wise semantic relevance at different velocity scales as shown in Fig. 1(b). With multi-velocity feature alignment, we can achieve more accurate matching to some extent in the collaboration of multiple-velocity features. However, with only limited support videos, the learned multi-velocity action features are not up to adhere to the underlying motion semantic without extra semantic constraints. Actually, the human perception system possesses a distinctive visual mechanism known as cognitive penetrability [18]. This mechanism employs prior linguistic knowledge to adjust visual perception processing towards semantic-relevant features. Hence, we argue that it is necessary to leverage text information to avoid multi-velocity features deviating from the underlying motion semantic for the rare support samples.
@@
Inspired by the above insights, we propose a novel Multi-Velocity Progressive-alignment (MVP-Shot) framework to learn and align multi-velocity action features in a progressive manner, as shown in Fig. 1(b). More specifically, we present a novel Progressive Semantic-Tailored Interaction (PSTI) module to capture multiple action features from different velocity scales, in which we recurrently inject velocity-tailored text information at different velocities to enhance velocity-aware semantic features. Thus rich semantic prior knowledge is fully utilized to avoid multiple velocity features deviating from the original motion semantic. To cooperate with the PSTI module, we further design a Multi-Velocity Feature Alignment (MVFA) module to capture pair-wise semantic similarity scores at different velocity scales, and then merge all similarity scores in a residual fashion to facilitate coherent velocity-aware matching between videos. In this way, our proposed MVP-shot combines the above two modules to compensate for each other, which enables more robust and more accurate few-shot matching.
@@
To comprehensively evaluate our method, we conduct experiments on four gold-standard datasets (i.e., HMDB51 [19], UCF101 [20], Kinetics [21], and SSv2-small [22]). We empirically prove that our approach surpasses all existing state-of-the-arts and yields solid performance gains, e.g., $3.2\%$ accuracy gains on HMDB51 dataset under the $5$-way $1$-shot setting. Furthermore, we perform thorough ablation studies to dissect each component, both quantitatively and qualitatively.
@@
  • We propose a novel MVP-Shot framework to enhance Few-Shot Action Recognition (FSAR) matching by learning and aligning reliable multi-velocity action features in a progressive way.
@@
  • To align multi-velocity action features covering different temporal dynamics, we present a new Multi-Velocity Feature Alignment (MVFA) module to measure the similarity of action features from different velocities in a residual fashion. To the best of our knowledge, this is the first work to exploit multi-velocity matching for FSAR.
The rest of this paper is organized as follows: Section §II describes the related works including few-shot image classification, few-shot action recognition, and transferring CLIP models for video recognition. Section §III presents the proposed MVP-shot in detail for few-shot action recognition. Section §IV reports the results and analysis of extensive experiments, followed by the conclusion in Section §V.
@@ II Related Work
In this section, we review several lines of researches closely related to the paper, including few-shot image classification (§II-A), few-shot action recognition (§II-B), and transferring CLIP models for video recognition (§II-C).
@@ II-A Few-Shot Image Classification
The few-shot learning (FSL) task, aiming to recognize new classes with limited samples, is an important research direction in computer vision. Existing FSL methods are roughly categorized into three groups: augmentation-based methods, optimization-based methods, and metric-based methods. Augmentation-based methods [23, 24, 25] usually learn to increase the number of training samples or features to augment model training and improve the generalization of few-shot models, mainly including hallucinating additional training examples  [23] and adversarial generation  [24]. Optimization-based methods [26, 27, 28, 29] learn a good initialization for the model or learn to update the parameters of the model properly so that the model can quickly adapt to new tasks with limited fine-tuning updates. Typical methods include MAML [26] and other variants [27, 28, 29]. Metric-based methods [30, 31, 32, 33, 34, 35, 36] leverage visual encoders to map images into a common embedding space and compare query and support images through different distance metrics, such as Euclidean distance [37, 30] and cosine similarity [38]. For example, Prototypical Network [39] can compute the Euclidean distances between the class prototypes of the query set and support set to obtain classification results. Our work also falls into the metric-based line and tries to solve the more challenging few-shot action recognition (FSAR) with diverse spatiotemporal dependencies. Our work focuses on making full use of temporal information to obtain and align semantic-related action features at multi-velocity levels.
Figure 2: The architecture of the proposed Multi-Velocity Progressive-Alignment (MVP-shot) Framework. Given the support and query videos, we first leverage a feature extractor to encode frame features. Then, based on frame features, we apply a Progressive Semantic-Tailored Interaction (PSTI) module to capture multi-velocity action features. Accordingly, we employ a Multi-Velocity Feature Alignment (MVFA) module to capture pair-wise semantic relevance at different velocity scales. For the convenience of illustration, other support videos in a few-shot task are not displayed in the figure.
@@ II-B Few-Shot Action Recognition
Few-shot action recognition aims to learn to categorize an unlabeled query video into one of the action categories in the support set with limited samples. Compared to images, video has richer spatio-temporal contextual information and temporal diversity. Therefore, it is not reasonable to directly apply few-shot image classification methods to the few-shot video recognition task. The existing few-shot action recognition methods [7, 40, 41, 42] mainly focus on metric-based learning to exploit the temporal cues. On the one hand, some approaches [43, 41] aim to enhance feature representation for few-shot action recognition. STRM [13] utilizes the spatiotemporal enrichment module, comprising a local patch-level enrichment sub-module and a global frame-level enrichment sub-module, to learn higher-order temporal representations. ITANet [44] introduces joint spatiotemporal modeling techniques to achieve implicit temporal alignment and matching. HyRSM [7] introduces a novel hybrid relation module to capture task-specific features, and SloshNet [42] designs a feature fusion architecture search module to automatically search for the best combination of the low-level and high-level spatial features. On the other hand, some methods focus on class prototype matching [11, 45, 46] strategies. Among them, OTAM [11] maintains the frame order in video and aligns the two video sequences using an improved dynamic time-warping algorithm. TARN [15] achieves temporal alignment by using an attention mechanism between the segment embeddings of the videos. TRX [14] matches each query sub-sequence with all sub-sequences in the support set and thus establishes query-specific video matching. In a word, the above methods focus on frame-level alignment or segment-level alignment solely. However, these single-scale approaches do not handle action instances with diverse speeds effectively. Unlike these previous methods, our work combines multi-velocity feature representation and multi-velocity feature alignment to compensate for each other, which enables more robust and more accurate few-shot matching.
@@ II-C Transferring CLIP Models for Video Recognition
With deep learning advancements [47, 48, 49, 50], significant progress has been made in the field of action recognition [51, 52, 53, 54] in recent years. Recently, image-language pretraining[55, 56] has achieved remarkable progress due to its powerful visual representations aligned with rich linguistic semantics. One of the most representative works is CLIP[56], which achieves impressive “zero-shot” transferability and generalization capacity. CLIP projects images and natural language descriptions to a common feature space through two separate encoders for contrastive learning. Thus these pre-trained models have been extended to various downstream tasks and achieve impressive performance, including image classification[57, 32] and video understanding[58, 59]. Inspired by these successes, CLIP-FSAR[60] firstly transfers CLIP for few-shot action recognition to integrate both visual features and auxiliary text features. However, CLIP models are image pre-trained models, while videos have richer spatio-temporal contextual information compared to images. Current methods only propose an additional temporal module to extend image pre-trained models, which are unable to take full advantage of temporal information in videos. To perform temporal modeling effectively, we propose a Semantic-Tailored Interaction module including two complementary mechanisms to inject semantic information into the visual features, which allow the interaction on the temporal and the channel dimensions, respectively. The Semantic-Tailored Interaction module fully utilizes semantic information of class names to obtain more discriminative video features and enhance the representation of video prototypes for few-shot learning.
@@ III Method
In this section, we first introduce the problem setup of the few-shot action recognition task (§III-A). Then, we present an architecture overview of the proposed MVP-shot (§III-B), as shown in Fig. 2. Finally, we describe each component of our proposed method in detail (§III-C and §III-D).
@@ III-A Problem Statement
The objective of few-shot action recognition is to learn a model that generalizes its performance to identify unseen categories using only a few samples. To make the training phase more faithful to the test phase, we adopt the episodic training manner [61] for few-shot adaptation as in previous work. In each episodic task, there are two sets, which consist of support set S and query set Q. The support set consists of N different action categories, and each category comprises K support videos, which typically range from $1$ to $5$. The problem setting like this is termed as the N-way K-shot problem. The goal is to correctly classify the sample in the query set into one of the N categories in the support set. Specifically, in the few-shot learning setting, the whole dataset is divided into a base set $\mathcal{D}_{base}=\{(v_{i},y_{i}),y_{i}\in\mathcal{C}_{base}\}$ and a novel set $\mathcal{D}_{novel}=\{(v_{i},y_{i}),y_{i}\in\mathcal{C}_{novel}\}$, wherein $y_{i}$ represents the action label of video $v_{i}$. Note that the categories of $\mathcal{D}_{base}$ and $\mathcal{D}_{novel}$ are sample-wise non-overlapping, $\mathcal{C}_{base}\cap\mathcal{C}_{novel}=\emptyset$. We utilize $\mathcal{D}_{base}$ to train our network for obtaining a robust and generalized model, and we use $\mathcal{D}_{novel}$ to evaluate the effectiveness of the proposed method.
@@ III-B Architecture Overview
The overall architecture of our proposed MVP-shot is illustrated in Figure 2. In a standard N-way K-shot episodic task, there is a support $S=\{s_{1},s_{2},...,s_{N\times K}\}$ consisting of $N$ action categories and $K$ videos per category,, where $s_{i}\in\mathbb{R}^{T\times 3\times H\times W}$ and $T$ is the number of sparsely sampled video frames. For the convenience of formulation, we consider the $N$-way $1$-shot task to present our framework, and the query set $Q$ contains a single video $q$. We first adopt the visual encoder of CLIP to generate the support features $F_{S}=\{f_{s_{1}},f_{s_{2}},...,f_{s_{N}}\}$ and query feature $f_{q}$, where $f_{i}=\{f^{1}_{i},f^{2}_{i},...,f^{T}_{i}\},f^{j}_{i}\in\mathbb{R}^{C}$ and $C$ is the channel number. We also use the text encoder of CLIP to obtain the textual embeddings of corresponding category natural language descriptions. Then we propose a Progressive Semantic-Tailored Interaction (PSTI) module to obtain semantic-related action features from different velocities. With such multi-velocity feature representation, we utilize a Multi-Velocity Feature Alignment (MVFA) module to measure the similarity of the specific semantic information at different temporal scales and produce the final predicted probabilities after a fusion operation. In our proposed MVP-shot, visual features from different velocity scales are complementary, and they work collaboratively to significantly improve classification performance.

### v4

@@ Abstract
Recent few-shot action recognition (FSAR) methods typically perform semantic matching on learned discriminative features to achieve promising performance. However, most FSAR methods focus on single-scale (e.g., frame-level, segment-level, etc.) feature alignment, which ignores that human actions with the same semantic may appear at different velocities. To this end, we develop a novel Multi-Velocity Progressive-alignment (MVP-Shot) framework to progressively learn and align semantic-related action features at multi-velocity levels. Concretely, a Multi-Velocity Feature Alignment (MVFA) module is designed to measure the similarity between features from support and query videos with different velocity scales and then merge all similarity scores in a residual fashion. To avoid the multiple velocity features deviating from the underlying motion semantic, our proposed Progressive Semantic-Tailored Interaction (PSTI) module injects velocity-tailored text information into the video feature via feature interaction on channel and temporal domains at different velocities. The above two modules compensate for each other to make more accurate query sample predictions under the few-shot settings. Experimental results show our method outperforms current state-of-the-art methods on multiple standard few-shot benchmarks (i.e., HMDB51, UCF101, Kinetics, and SSv2-small).
@@
Few-shot learning, Action recognition, Few-shot action recognition, Mutli-velocity feature matching
@@ I Introduction
Due to the increasing demand for video analysis in real-world scenarios, significant progress has been made in action recognition [1, 2, 3, 4, 5] recently. Despite this, this success still relies heavily on extensive manually-labeled samples, and collecting these samples is expensive and labor-intensive. This significant bottleneck severely constrains the broader applicability of this task. To alleviate the data-hunger issue, few-shot action recognition (FSAR) is a promising direction and has gained widespread attention [6, 7, 8, 9, 10] in recent years. The objective of FSAR is to identify unseen action categories with only a few labeled video examples.
@@
Current mainstream FSAR approaches are built upon the metric-based meta-learning paradigm, which learns discriminative class prototype representation and predicts action category labels based on learned distance metrics. Most metric-based methods perform frame-level alignment [11, 7, 12, 13, 9] on learned discriminative frame-level features, which measure the similarity between different videos via diverse matching strategies [11, 7, 14]. In addition, some methods perform segment-level alignment [15, 16, 17] on obtained segment-level features. Typically, TARN [15] first uniformly samples segments from videos, and performs segment-level alignment based on segment features extracted by 3D-CNN. Although existing FSAR methods have achieved remarkable performance, these methods still mainly focus on single-scale feature alignment (i.e., frame-level alignment and segment-level alignment), as shown in Fig. 1(a), which is insufficient to reflect the diversity of action speeds. Therefore, these single-scale approaches often result in inaccurate matching when dealing with action instances with diverse speeds. In real-world scenarios, different individuals may perform the same action class at different velocities. Take the action “ride horse” as an example, different persons act at different velocities. Specifically from a visual perspective, certain horses may be perceived sprinting at a rapid pace, while others may be observed moving at a more leisurely walk.
@@
In our paper, we focus on multi-velocity feature alignment in FSAR, where multi-velocity features are utilized to capture pair-wise semantic relevance at different velocity scales. With multi-velocity feature alignment, we can achieve more accurate matching to some extent in the collaboration of multiple-velocity features. However, with only limited support videos, the learned multi-velocity action features are not up to adhere to the underlying motion semantic without extra semantic constraints. Actually, the human perception system possesses a distinctive visual mechanism known as cognitive penetrability [18]. This mechanism employs prior linguistic knowledge to adjust visual perception processing towards semantic-relevant features. Hence, we argue that it is necessary to leverage text information to avoid multi-velocity features deviating from the underlying motion semantic for the rare support samples.
@@
In light of the above insights, we propose a novel Multi-Velocity Progressive-alignment (MVP-Shot) framework to learn and align multi-velocity action features in a progressive manner (Fig. 1(b)). More specifically, we present a novel Progressive Semantic-Tailored Interaction (PSTI) module to capture multiple action features from different velocity scales, in which we recurrently inject velocity-tailored text information at different velocities to enhance velocity-aware semantic features. Thus rich semantic prior knowledge is fully utilized to avoid multiple velocity features deviating from the original motion semantic. To cooperate with the PSTI module, we further design a Multi-Velocity Feature Alignment (MVFA) module to capture pair-wise semantic similarity scores at different velocity scales, and then merge all similarity scores in a residual fashion to facilitate coherent velocity-aware matching between different videos. In this way, our proposed MVP-shot combines the above two modules to compensate for each other, which enables more robust and more accurate few-shot feature matching.
@@
We conduct experiments on four gold-standard datasets (i.e., HMDB51 [19], UCF101 [20], Kinetics [21], and SSv2-small [22]) to comprehensively evaluate the effectiveness of our MVP-shot. We empirically prove that our MVP-shot surpasses all existing state-of-the-arts and yields solid performance gains, e.g., $3.2\%$ accuracy gains on HMDB51 dataset on the $5$-way $1$-shot setting. Furthermore, we perform thorough ablation studies to dissect each component, both quantitatively and qualitatively.
@@
  • We propose a novel MVP-Shot framework to conduct Few-Shot Action Recognition (FSAR) matching by learning and aligning reliable multi-velocity action features in a progressive way.
@@
  • To align multi-velocity action features covering different temporal dynamics, we present a new Multi-Velocity Feature Alignment (MVFA) module to measure the similarity of action features from different velocities in a residual fashion. To the best of our knowledge, this is the pioneering effort to exploit multi-velocity matching for FSAR.
We organize the rest of this paper as follows: Section §II describes several lines of related works regarding few-shot image classification, few-shot action recognition, and transferring CLIP models for video recognition. Section §III presents the proposed MVP-shot in detail for FSAR. Section §IV reports comparison results and analysis of extensive experiments. Finally, the conclusion is provided in Section §V.
@@ II Related Work
In this section, we introduce several lines of research closely related to our MVP-shot, regarding few-shot image classification (§II-A), few-shot action recognition (§II-B), and transferring CLIP models for video recognition (§II-C).
@@ II-A Few-Shot Image Classification
The few-shot learning (FSL) task aims to recognize unseen classes with limited labeled samples. Mainstream FSL methods are roughly categorized into three groups: i) Augmentation-based methods [23, 24, 25] usually learn to increase the number of features or training samples to augment model training and improve few-shot model’s generalization ability, mainly including hallucinating additional training examples  [23] and adversarial generation  [24]; ii) Optimization-based methods [26, 27, 28, 29] learn an effective model initialization state or develop techniques to properly update model parameters, enabling rapid adaptation to new tasks with minimal fine-tuning steps for enhanced generalization.Typical methods include MAML [26] and other variants [27, 28, 29]; iii) Metric-based methods [30, 31, 32, 33, 34, 35, 36] leverage visual encoders to map images into a common embedding space and predict query (i.e., test sample) labels with respect to different distance metrics, such as Euclidean distance [37, 30] and cosine similarity [38]. For example, Prototypical Network [39] measures the similarity of test samples and the class prototypes of the support set by Euclidean distance metric and makes few-shot predictions based on obtained similarity scores. Our work belongs to the metric-based methods and tries to solve challenging few-shot action recognition (FSAR), which aims to identify unlabeled videos with complex spatiotemporal structures using only a few video examples. Our work focuses on making full use of temporal information to obtain and align semantic-related action features at multi-velocity levels.
Figure 2: The overview of the proposed Multi-Velocity Progressive-Alignment (MVP-shot) Framework. The support and query videos are first fed into a visual encoder to extract frame-level features. Then, based on frame features, we apply a Progressive Semantic-Tailored Interaction (PSTI) module to capture multi-velocity action features. Accordingly, we employ a Multi-Velocity Feature Alignment (MVFA) module to measure pair-wise semantic similarity at different velocity scales. For clarity, the figure does not depict other support videos in the few-shot task.
@@ II-B Few-Shot Action Recognition
Few-shot action recognition (FSAR) learns discriminative class prototype representation and categorizes unlabeled query video samples with a few labeled samples per class (i.e., support samples) by performing prototype-query matching based on distance metrics. Different from images, video has temporal diversity and richer spatio-temporal contextual information. Therefore, directly applying few-shot image classification methods to few-shot video recognition tasks is not appropriate. The existing few-shot action recognition methods [7, 40, 41, 42] mainly follow a metric-based learning paradigm to exploit the temporal cues. As the primary step, some approaches [43, 41] aim to learn robust feature representation for few-shot action recognition. STRM [13] performs spatial feature interaction in each frame and temporal feature interaction by attention mechanisms in a cascaded manner, so as to learn discriminative feature representations. SloshNet [42] designs an architecture search module for feature fusion, aiming to automatically discover the optimal integration of spatial features from different levels. ITANet [44] introduces joint spatiotemporal modeling techniques to achieve implicit temporal alignment and matching. HyRSM [7] makes full use of all videos in a meta-learning task to learn task-specific patterns by hybrid relation module. Other methods attempt to design robust class prototype matching [11, 45, 46] strategies. Among them, OTAM [11] maintains the frame order in video and aligns the two video sequences using an improved dynamic time-warping algorithm. TARN [15] utilizes an attention mechanism between segment visual features to achieve temporal alignment. TRX [14] establishes query-specific video matching by matching sub-sequences between different videos. In a word, the aforementioned methods focus on segment-level alignment or frame-level alignment solely. However, these single-scale approaches do not handle action instances with diverse speeds effectively. Unlike these previous methods, our work combines multi-velocity feature representation and multi-velocity feature alignment to compensate for each other, which enables more robust and more accurate few-shot matching.
@@ II-C Transferring CLIP Models for Video Recognition
With deep learning advancements [47, 48, 49, 50], significant progress has been made in the field of action recognition [51, 52, 53, 54] in recent years. Recently, image-language pretraining[55, 56] has achieved remarkable progress due to its powerful visual representations aligned with rich linguistic semantics. CLIP[56] is one of the most representative works, which achieves impressive “zero-shot” transferability and generalization capacity. CLIP is a vision-language transformer capable of jointly understanding both images and text by contrastive learning. Various downstream tasks begin to widely adopt these pre-trained models, showcasing impressive performance across the board, including image classification[57, 32] and video understanding[58, 59]. Inspired by these successes, CLIP-FSAR[60] firstly makes use of CLIP to integrate auxiliary semantic information (provided by class names) and visual features for FSAR. However, CLIP models are trained on large-scale image pairs, and different from images, videos contain richer spatiotemporal contextual information. Current methods only propose an additional temporal module to extend image pre-trained models, which are unable to take full advantage of temporal information in videos. To perform temporal modeling effectively, we propose a Semantic-Tailored Interaction module including two complementary sub-modules to inject semantic information provided by class names into the visual features, which performs feature interaction on the temporal and the channel dimensions. The Semantic-Tailored Interaction module fully utilizes semantic information provided by class names to obtain more discriminative video features and enhance video class prototype representation for few-shot learning.
@@ III Method
In this section, we first introduce the problem statement of the few-shot action recognition task (§III-A). Then, as shown in Fig. 2, an architecture overview of our proposed MVP-shot is presented (§III-B). Finally, we describe in detail each component of our proposed MVP-shot (§III-C and §III-D).
@@ III-A Problem Statement
The goal of few-shot action recognition is to learn a model that has a strong generalization capability to identify unseen categories given a few (e.g., one or five) labeled samples per class. In the few-shot learning setting, the whole dataset is divided into a training set $\mathcal{D}_{base}=\{(v_{i},y_{i}),y_{i}\in\mathcal{C}_{base}\}$ and a testing set $\mathcal{D}_{novel}=\{(v_{i},y_{i}),y_{i}\in\mathcal{C}_{novel}\}$, where $y_{i}$ denotes the action label of video $v_{i}$. Note that the categories of $\mathcal{D}_{base}$ and $\mathcal{D}_{novel}$ are sample-wise non-overlapping, i.e., $\mathcal{C}_{base}\cap\mathcal{C}_{novel}=\emptyset$. Under the few-shot setting, we follow the episodic training paradigm [61] to train a meta-learning model for FSAR. In each episodic task (i.e., a standard N-way K-shot task), there are two sets (i.e., query set Q and support set S) sampling N different action categories from training set. The support set consists of K support videos per category, which typically range from $1$ to $5$, and the query set consists of a fraction of the rest videos from N action categories. The goal is to correctly classify the query sample into one of the N categories in the support set. During the inference phase, we randomly sample multiple episodic tasks from $\mathcal{D}_{novel}$ to evaluate the model’s generalization capabilities.
@@ III-B Architecture Overview
We illustrate the overall architecture of our proposed MVP-shot in Figure 2. In a standard N-way K-shot episodic task, the support set $S=\{s_{1},s_{2},...,s_{N\times K}\}$ consists of $N$ action categories and $K$ videos per category, where $s_{i}\in\mathbb{R}^{T\times 3\times H\times W}$ and $T$ is the number of sparsely sampled video frames. For the sake of clarity, we present our framework using $N$-way $1$-shot task, and a single video $q$ is included in the query set $Q$. The visual encoder of CLIP first processes video inputs to generate query video feature $f_{q}$ and the support video features $F_{S}=\{f_{s_{1}},f_{s_{2}},...,f_{s_{N}}\}$, where $f_{i}=\{f^{1}_{i},f^{2}_{i},...,f^{T}_{i}\},f^{j}_{i}\in\mathbb{R}^{C}$ and $C$ denotes the channel number. In addition, the corresponding category names are input into the text encoder of CLIP to obtain textual features. Then we propose a Progressive Semantic-Tailored Interaction (PSTI) module to obtain semantic-related action features from different velocities. With such multi-velocity feature representation, we utilize a Multi-Velocity Feature Alignment (MVFA) module to measure the similarity of the specific semantic information at different temporal scales and produce the final predicted probabilities after a fusion operation. In our proposed MVP-shot, visual features from different velocity scales are complementary, and they work together to enable more robust and more accurate few-shot feature matching.